OpenStudy (anonymous):

How to simplify 3^(1/n) * 4^(1/n) ?

geerky42 (geerky42):

Like that? \[\Large 3^{\frac{1}{n}} \cdot 4^{\frac{1}{n}}\]

OpenStudy (anonymous):

yes

geerky42 (geerky42):

In that case, since both have same exponent, we can do that: \[\Large 3^{\frac{1}{n}} \cdot 4^{\frac{1}{n}} \Rightarrow (3 \cdot 4)^{\frac{1}{n}}\]